EDITORIAL

Terry Goodman

We at MJMS hope your school year has started off well. The beginning of a new year is always exciting and I am sure you are already busy as you encounter new challenges.

Here at Central we are beginning to incorporate the graphing calculator into several of our courses. Since we are in the early stage in the use of this technology, we find we are using some trial and error in our experimentation.

In some classes the calculator is used by the instructor as a demonstration tool, while in other classes each student rents a calculator to use throughout the course. Our initial efforts indicate that the graphing calculator holds great promise as a powerful tool for mathematics students.

In using the calculator with a calculus class I am teaching, I have found that the students took some time to become comfortable with this tool. Once they began to use the calculator, however, their enthusiasm increased and they were willing to investigate a wider variety of problems than my previous classes had been.

We have been able to more fully consider geometric interpretation of topics such as solving polynomial equations, functions and their inverses, points of intersection, rational functions, limits, continuity, and removable and non-removable discontinuities. Students are able to even explore problems that do not have only “nice” numbers in them. Graphing a function such as

f(x) = 25(2[[\frac{t + 2}{2}]]), t ≥ 0

by hand can be quite a chore. The graphing calculator allows students to readily investigate the behavior of this function.

Since we are just beginning to use graphing calculators, we are seeking information from others who are using this technology. MJMS will welcome manuscripts describing uses of graphing calculators and other technologies. We all benefit from the sharing of ideas as each individual strengthens his/her understanding. Further, it may be possible for us to help our students better understand and apply mathematics by using powerful tools such as the graphing calculator.

Send us your ideas. We want to share them with others.

We would also like to acknowledge and thank the following people who served as referees for MJMS during this past year.

List of Referees

Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville John E. Bernard, Pan American University Albert Dixon, College of the Ozarks John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Russell Euler, Northwest Missouri State University Joe Flowers, Northeast Missouri State University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Leon Hall, University of Missouri–Rolla Dennis R. Harmon, Missouri Southern State College Larry K. Johnson, Central Missouri State University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Kandasamy Muthvel, University of Wisconsin–Oshkosh Krzysztof Ostaszewski, University of Louisville James W. Petticrew, University of Texas–Pan American Bob Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Jau–Shyong Shiue, University of Nevada–Las Vagas Shing S. So, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Alvin Tinsley, Central Missouri State University Bobby Neal Winters, Pittsburg State University